OpenStudy (anonymous):

A teacher grades homework by randomly selecting 4 out of 10 homework problems to grade. How many different groups of 4 problems can be taken from the assignment of 10 problems?

OpenStudy (lgbasallote):

use combinations \[nCr = \frac{n!}{r!(n-r)!}\] so \[10C4 = \frac{10!}{4!(10-4)!}\] does that help?

OpenStudy (anonymous):

so do i multiply

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (anonymous):

how do i get the different groups

OpenStudy (lgbasallote):

you solve what i wrote above \[\frac{10!}{4!(10-4)!}\] you know what those notations mean right?

OpenStudy (anonymous):

36

OpenStudy (lgbasallote):

that's not right....what did you do?

OpenStudy (anonymous):

4x10=-4

OpenStudy (lgbasallote):

4x10 is not -4.... what do you mean 4x10 = -4?

OpenStudy (anonymous):

4x10=40-4=36

OpenStudy (lgbasallote):

hmm well no...that's not how to do it..

OpenStudy (lgbasallote):

10! = 10 x 9 x 8 x 7 x 6 x 5 x 4 x 3 x 2 x 1 4! = 4 x 3 x 2 x 1 (10-4)! = 6! = 6 x 5 x 4 x 3 x 2 x 1 now can you solve it?

OpenStudy (anonymous):

3628800 24 720

OpenStudy (lgbasallote):

??

OpenStudy (anonymous):

i just multiply

OpenStudy (anonymous):

so it 4

OpenStudy (anonymous):

?